A positive point charge (q = +7.55 x 10-8 C) is surrounded by an equipotential surface A, which has a radius of rA = 2.99 m. A positive test charge (q0 = +3.53 x 10-11 C) moves from surface A to another equipotential surface B, which has a radius rB. The work done by the electric force as the test charge moves from surface A to surface B is WAB = -9.38 x 10-9 J. Find rB.
Answer #1

\tiny Work..done..by..electric..force = \int \overrightarrow{F}.\overrightarrow{dx}=-\Delta PE= Kqq0 / r_{a} - Kqq0 / r_{b}

=> -9.38 x 10-9 J = Kqq0/ ra - Kqq0/ rb

=> -9.38 x 10-9 J = 9 x 109 x 7.55 x 10-8 x3.53 x 10-11 x(1/ 2.99 -   1/ rb)

=> -0.391 = (1/ 2.99 -   1/ rb)  

=>    1/ rb = 0.73

=> rb = 1.38 m